### v1

# FPGA-tailored algorithms for real-time decoding of quantum LDPC codes

Abstract
Real-time decoding is crucial for fault-tolerant quantum computing but likely requires specialized hardware such as field-programmable gate arrays (FPGAs), whose parallelism can alter relative algorithmic performance. We analyze FPGA-tailored versions of three decoder classes for quantum low-density parity-check (qLDPC) codes: message passing, ordered statistics, and clustering. For message passing, we analyze the recently introduced Relay decoder and its FPGA implementation; for ordered statistics decoding (OSD), we introduce a filtered variant that concentrates computation on high-likelihood fault locations; and for clustering, we design an FPGA-adapted generalized union-find decoder. We design a systolic algorithm for Gaussian elimination on rank-deficient systems that runs in linear parallel time, enabling fast validity checks and local corrections in clustering and eliminating costly full-rank inversion in filtered-OSD. Despite these improvements, both remain far slower and less accurate than Relay, suggesting message passing is the most viable route to real-time qLDPC decoding.
1 Introduction and summary of main results
Fault-tolerant quantum computation (FTQC) enables large-scale quantum algorithms by using quantum error correction (QEC) to protect logical information from the inevitable noise in physical qubits and operations. A central challenge in realizing FTQC is the implementation of a real-time decoder, a classical algorithm that corrects errors fast enough to keep pace with successive QEC cycles. For superconducting qubits, with microsecond QEC cycle times [Battistel_2023], this will likely require specialized computing hardware such as field-programmable gate arrays (FPGAs) rather than central processing units (CPUs), which in turn influences the choice of decoder because the platforms differ fundamentally. Unlike CPUs, which are the target platforms for most decoders, FPGAs comprise millions of small, interconnected processing elements pre-configured to execute fixed local functions in parallel to collectively realize a specified computation.
The appropriate choice of decoder can also depend on the underlying QEC code. Recent theoretical advances in quantum low-density parity-check (qLDPC) codes, combined with improving qubit connectivity, have made these codes increasingly attractive alternatives to the surface code, offering comparable protection with far fewer physical qubits. However, most existing FPGA decoders [Battistel_2023, vittal2023astrea, wu2025micro, ziad2024local, liyanage2023scalable] exploit specific features of the surface code that do not apply to qLDPC codes. This motivates a comparative study of decoder classes suitable for real-time FPGA implementation of general qLDPC codes.
Figure 1: Classes of qLDPC decoder. Among these, classes (iv)–(vi) are identified as the most suitable for real-time decoding in FPGAs and are evaluated in this work.
Existing decoders for general qLDPC codes can be grouped into six classes (Figure 1): (i) SAT-reduction, (ii) search-based, (iii) neural-network, (iv) message-passing, (v) ordered-statistics, and (vi) clustering. SAT-reduction/SMT-based [Berent_2024, noormandipour2024maxsatdecodersarbitrarycss, shutty2022decoding, Zhou2025] and search-based [Ott2025, beni2025tesseract, wu2025minimum] decoders have variants that guarantee minimum-weight corrections, but they remain orders of magnitude slower than other approaches and rely on complex data structures and conditional logic that are inefficient on FPGAs [besta2019graphprocessingfpgastaxonomy, Halstead2011, Luo2017]. Neural-network decoders [hu2025efficient, blue2025machine, Gong2024, Ninkovic2024, maan2025decoding] appear too slow for microsecond QEC cycles, and are better suited to GPUs than to FPGAs. We therefore focus on the remaining three classes: message passing, ordered-statistics decoding (OSD), and clustering. Our primary technical contributions are:
  1. 1.
    An FPGA-tailored filtered-OSD, avoiding standard OSD’s costly explicit matrix inversion.
  2. 2.
    An FPGA-tailored cluster decoder based on a generalized Union–Find decoder [delfosse2022toward].
  3. 3.
    A systolic-array solver for rank-deficient linear systems, used in both of the above decoders.
  4. 4.
    A comparison of logical error rates for these decoders against Relay message-passing [muller2025improved] under fixed FPGA cycle budgets to identify the most promising class for real-time decoding.
Our FPGA-tailored algorithms use basic operations organized for efficient FPGA execution, yielding runtime and resource costs that differ sharply from CPU implementations of the same logic. We model decoding via a binary decoding matrix $H\in\mathbb{F}_{2}^{M\times N}$ that maps fault vectors $x\in\mathbb{F}_{2}^{N}$ to syndromes $\sigma\in\mathbb{F}_{2}^{M}$ through $Hx=\sigma$. Given $\sigma$, the decoder seeks a likely correction $y$ with $Hy=\sigma$.
Message-passing decoders.
These algorithms exchange information iteratively between check and fault nodes in a graphical representation of the decoding matrix, with local update rules that map naturally onto parallel FPGA architectures [pamuk2011fpga, valls2021syndrome]. The standard example is Min-Sum belief propagation (BP), which often fails to converge for quantum LDPC codes [poulin2008on, maan2025decoding] in part due to loops and symmetries in the decoding graph. Recent advances address these issues through decimation to break loops [du2022stabilizer, gong2024toward, yao2024belief, tsubouchi2025degeneracy, du2024check, yin2024symbreak, demarti2024closed, demarti2024almost], incorporating memory to damp message oscillations [chytas2025enhanced, murphy2013loopy, muller2025improved], and decoder ensembling [koutsioumpas2025automorphism], while others learn modified message-update rules by pretraining [liu2019neural]. The recently introduced Relay decoder [muller2025improved] performs a sequence of BP runs (“legs”) with varying memory-strength parameters, progressively breaking trapping sets to improve convergence and accuracy. We use Relay as a representative benchmark for the message-passing class, drawing on published FPGA performance analyses [muller2025improved, maurer2025realtimedecoding].
Ordered-statistics decoders.
OSD algorithms [panteleev2021degenerate, fossorier1995soft, fossorier2002iterative] leverage an unconverged pre-decoder’s output by reprocessing bits according to their estimated reliability. Following Ref. [roffe2020], the standard OSD selects a set $S$ of linearly independent columns of $H$, ordered by likelihood, with $|S|=\operatorname{rank}(H)$. The corresponding submatrix $H_{S}\in\mathbb{F}_{2}^{M\times|S|}$ is inverted to obtain a correction $x=H_{S}^{-1}\sigma$ supported on $S$. However, standard OSD is impractical on FPGAs [valls2021syndrome] because detecting independence and inverting the large $\operatorname{rank}(H)\!\times\!\operatorname{rank}(H)$ matrix dominate runtime, even when optimized [bascones2025exploring].
Our first technical contribution is an FPGA-tailored variant, filtered-OSD, which avoids explicit matrix inversion by solving $H_{S^{\prime}}x=\sigma$ directly on a smaller, potentially rank-deficient submatrix $H_{S^{\prime}}$ [panteleev2021degenerate]. We obtain $S^{\prime}$ from a pre-decoder by discarding low-probability columns. Because $|S^{\prime}|\ll\operatorname{rank}(H)$, column sorting and linear solving are both substantially accelerated. Sparsity further reduces cost because many rows of $H_{S^{\prime}}$ are zero and can be skipped by the solver. This approach requires an FPGA-tailored solver for rank-deficient binary systems, described next.
Systolic rank-deficient solver.
We provide a new linear-time systolic-array algorithm that solves $Ax=b$ for any matrix $A\in\mathbb{F}_{2}^{m\times n}$ and vector $b\in\mathbb{F}_{2}^{m}$. Systolic arrays are a class of highly parallel algorithms that implement computations on grids of processing elements and pass data locally each cycle, making them particularly well suited to FPGAs. Existing FPGA solvers typically assume that $A$ has full rank [gentleman1982matrix, Hochet1989, cosnard1986matching, robert1985resolution, wang1993systolic, parkinson1984compact, Rupp2006, rupp2011hardware, jasinski2010gf2, scholl2013hardware, scholl2014hardware, Shoufan2010, wang2017fpga, bascones2025exploring], while the only prior solver for arbitrary $A$ [hu2024universal] employs relatively complex processing elements. Our solver is simpler and may be of independent interest; it forms the computational core of our filtered-OSD and cluster decoder.
Cluster decoders.
This class of decoders [kovalev2013fault, delfosse2021almost, delfosse2022toward, higgott2023improved, hillmann2024localized, wolanski2024ambiguity] group nearby detection events in the decoding graph that are likely caused by the same underlying faults, iteratively growing and merging clusters until each can be locally corrected. Mathematically, by the end of the algorithm each cluster $C$ corresponds to a set of columns of $H$ with local corrections $x_{C}$ satisfying $H_{C}x_{C}=\sigma_{C}$, where $\sigma_{C}$ is the restriction of $\sigma$ to the neighborhood of $C$. The local corrections are then combined to obtain a global $x$ obeying $Hx=\sigma$. Initially, clusters may have no solution to $H_{C}x_{C}=\sigma_{C}$, indicating that they must be grown before becoming solvable. Cluster-decoder variants differ in their growth and merge rules and may exploit code structure. For example, the Union–Find decoder [delfosse2021almost] avoids Gaussian elimination by leveraging surface-code structure.
We introduce an FPGA-tailored version of the algorithm in Ref. [delfosse2022toward], which generalizes the Union–Find decoder [delfosse2017almost] to arbitrary qLDPC codes. This decoder uses the systolic solver to test cluster solvability and extract local corrections, and incorporates hardware-specific optimizations to maximize throughput. As in filtered-OSD, it uses a pre-decoder: faults with high likelihoods are accepted immediately [caune2023belief], while intermediate-likelihood faults are treated as erasures that seed the initial clusters together with unsatisfied syndrome nodes.
Figure 2: Cutoff-time performance curves for FPGA-tailored decoders. Logical error rate versus FPGA cycle budget for Relay, filtered-OSD, cluster decoding, and standard OSD [bascones2025exploring], with runs exceeding the cycle cutoff counted as failures. Relay achieves the lowest error rates across all tested budgets. Filtered-OSD starts up with significantly fewer cycles than standard OSD.
Real-time decoding analysis.
We compare the real-time feasibility of the aforementioned decoders in Figure 2 using Monte Carlo simulations of the gross and two-gross qLDPC codes [bravyi2024high, yoder2025tour] to estimate logical error rates within a fixed FPGA cycle budget and to assess resource requirements on a representative commercial FPGA device (see Section LABEL:sec:results). For a consistent comparison, we use the first leg of Relay as the pre-decoder for standard OSD, filtered-OSD, and cluster decoding, and all performance data in Figure 2 are reported starting from the end of that leg.
Both our FPGA-tailored decoders achieve substantial runtime reductions relative to the standard OSD implementation of Ref. [bascones2025exploring]. However, Relay reduces the logical error rate more rapidly with increasing cycle budget than the other three decoders, quickly reaching an order of magnitude lower rates for the gross code and two orders lower for the two-gross code. Although Figure 2 shows that Relay exhibits a long-tailed latency distribution, in Section LABEL:sec:real-time-decoding we show that this tail does not lead to problematic backlog for our decoding examples, as they satisfy the latency-tail conditions we derive in Section 2.2. These findings suggest that message-passing approaches, such as Relay, currently offer the most practical route to realizing real-time fault-tolerant decoding of large-scale quantum LDPC codes in FPGAs.
This overall conclusion appears unlikely to change without substantial new algorithmic advances in OSD or cluster decoding algorithms. Although our cycle counts and resource estimates are obtained at the architectural level rather than from full Verilog implementations (see Section LABEL:sec:results), the observed performance gaps are simply too large to be bridged by modest clock-frequency variations. Here is a clearer and slightly tighter version with no added length: OSD could improve logical error rates by adding a combination-sweep (CS) stage [roffe2020], but not enough to rival Relay [muller2025improved]. Moreover, although CS adds negligible overhead on CPUs (where OSD’s elimination cost dominates over the matrix–vector multiplications in CS) it would cause a substantial slowdown on FPGAs, where elimination and matrix–vector multiplication have comparable linear parallel-time cost. For similar reasons, although clustering could reuse partial elimination results [hillmann2024localized] for a speedup on CPUs, we expect little advantage on FPGAs where elimination is less of a bottleneck. Cluster decoders might also benefit from non-uniform, pre-decoder-biased growth [wolanski2024ambiguity], but such extensions would introduce conditional logic and data-management overheads that reduce throughput.
Outline.
In the remainder of this paper, we first provide background and definitions for QEC decoding and FPGA algorithms in Section 2 and Section LABEL:sec:fpga-background. Given the growing interest in using FPGAs for real-time decoding we take the opportunity to provide relatively self-contained technical introductions aimed at bridging the gaps between the communities. We then present our systolic solver in Section LABEL:sec:gateware_lse before presenting our FPGA-tailored decoders in Section LABEL:sec:fpga-osd and Section LABEL:sec:fpga-cluster, giving results and conclusions in Section LABEL:sec:results.
2 QEC decoding background and definitions
This section provides a brief overview of quantum error correction (QEC), which protects quantum information by redundantly encoding logical qubits into many physical qubits and continuously monitoring for errors. In this work, we consider a scenario where a dedicated circuit repeatedly measures the parity checks of a QEC code, producing classical syndrome data that signals the occurrence of faults. A decoder is a classical algorithm that processes the syndrome stream to determine the required corrections. For large-scale quantum computation, the decoder must operate in real time, processing syndrome data at the rate it is generated to prevent both stalling the quantum program and the accumulation of an unmanageable backlog [terhal2015quantum]. This work focuses on implementing such real-time decoders in classical hardware.
In the rest of this section, we first review QEC codes, circuits, and the decoding process in Section 2.1, followed by a discussion of the specific decoding algorithms relevant to this study.
2.1 QEC codes, circuits, noise and decoding
Codes
The most common class of quantum codes are stabilizer codes. A stabilizer code encodes $k$ logical qubits into $n$ physical qubits by defining the codespace as the joint +1+1 eigenspace of an abelian subgroup $S$ of the $n$-qubit Pauli group $\mathcal{P}_{n}$ that does not contain I-I. The subgroup $S$ is generated by nkn-k independent stabilizer generators, and its full group has size $2^{n-k}$. Logical operators correspond to Pauli operators in the normalizer $N(S)$ that commute with all elements of $S$ but are not in $S$ (up to a phase). The distance $d$ of the code is the minimum weight (number of qubits acted upon) of such a logical operator and quantifies the number of errors the code can reliably correct. We show an example in Figure 3.
In the simplest noise model, an unknown Pauli error $E$ acts on the system, after which the stabilizer generators are measured to produce a binary syndrome vector whose $1$ entries correspond to 1-1 measurement outcomes. Only generators that anticommute with $E$ produce 1-1 outcomes, allowing the syndrome to identify an error class. A decoder maps the syndrome to a Pauli correction $C$ with the same syndrome, returning the state to the codespace. Up to a phase, the net operator $CE$ either belongs to $S$, in which case error correction succeeds, or to $N(S)\setminus S$, in which case a logical error has occurred. Any error of weight strictly less than $d/2$ can be corrected.
[…]
Circuits
In practice, stabilizer generators cannot be measured directly because quantum hardware only supports a limited set of physical operations. Typical gate sets include single-qubit and two-qubit gates along with single-qubit measurements, and all stabilizer measurements must ultimately be decomposed into sequences of these elementary operations. A standard approach is to introduce ancilla qubits that interact with the data qubits via two-qubit gates so that measuring the ancilla indirectly measures the desired stabilizer generator (see Figure 3).
Physical operations are subject to noise, which is more accurately modeled at the circuit level. Because these noisy circuits produce unreliable stabilizer measurement outcomes, a standard strategy is to repeatedly execute the stabilizer measurement circuit. Syndrome bits are obtained by taking the parity of consecutive outcomes for each generator, detecting faults that affect either the data qubits or the measurement process itself.
Noise and decoding
In this work, we adopt the standard circuit noise model, in which each gate, measurement, or idle period (a qubit not involved in any operation) has the same duration and fails independently with probability $p$. A fault is modeled as the ideal operation followed by a random non-trivial Pauli error on its qubits, and in the case of a measurement, a flip of the reported outcome.
The complete mathematical description of the circuit decoding problem is specified by: a binary decoding matrix $H\in\mathbb{F}_{2}^{M\times N}$, a binary logical action matrix $A\in\mathbb{F}_{2}^{K\times N}$, and a fault probability vector $\mathbf{p}\in[0,1]^{N}$. There are $M$ detectors, which each take a value 0 or $1$ during a noisy run of the circuit, forming a bitstring called the syndrome, and $N$ distinct faults which occur independently, where the $j$th fault occurs with probability $p_{j}$. When the $j$th fault alone occurs, the observed syndrome is the $j$th column of $H$. When a set of faults occur together, their effect on the detectors is linear over $\mathbb{F}_{2}$. Let the bitstring $F\in\mathbb{F}_{2}^{N}$ specify a set of faults (where $F_{j}=1$ if and only if the $j$th fault occurs). The probability of $F$ is
$\Pr(F)=\prod_{j=1}^{N}(1-p_{j})\left(\frac{p_{j}}{1-p_{j}}\right)^{F_{j}}.$
The syndrome $\sigma(F)$ is then $\sigma(F)=HF$, where arithmetic (throughout this work, unless explicitly stated) is performed over $\mathbb{F}_{2}$. The logical action of $F$ is $a(F)=AF$.
A run of the noisy error correction circuit is modeled as follows: The circuit is executed, and a set of (unknown) faults $F$ occurs with probability $\Pr(F)$. The syndrome bitstring $\sigma(F)=HF$ is known, even though $F$ is not. The syndrome is passed to a decoder $\mathcal{D}$, which is a classical algorithm with knowledge of $H$, AA, and $\mathbf{p}$. The decoder outputs a correction $\hat{F}=\mathcal{D}(\sigma(F))$. If $a(F)+a(\hat{F})=A(F+\hat{F})=0$, then we say that the decoder succeeds; otherwise, it fails. Note that it is not necessary that $\hat{F}=F$ in order for the decoder to succeed.
We have not specified here exactly how to construct the decoding objects $H$, AA and $p$ from a specific circuit, or how to extract the syndrome $\sigma$ from measurement outcomes of the circuit in a given run, since the details of how this is done depends on the noise model. We address this in more detail in our paragraph on decoding examples at the end of this subsection.
Stochastic and erasure faults
It can be useful to separate faults into two classes: stochastic faults (which occur with small probability in unknown locations) and erasure faults (which occur with high probability, but in known locations). In practice, these can arise due to quite different physical processes, with stochastic faults arising, for example, from dephasing errors, and erasure faults arising, for example, from a qubit that goes missing (with our knowledge) and is then replaced. In the mathematical model we have presented, these two scenarios are simply covered by the probability of each fault — with the probability $p_{j}$ of a stochastic fault being some small value, and the probability of an erasure fault being $p_{j}=1/2$.
We can define a Tanner graph $\mathcal{G}$ for the decoding problem. $\mathcal{G}$ is a bipartite graph where each column of $H$ corresponds to a circular fault node, and each row corresponds to a square check node, and there is an edge connecting a pair of nodes if, and only if, the corresponding entry of $H$ is $1$. Given the graph $\mathcal{G}$, we define the nodes adjacent to node $j$ as $\mathcal{N}(j)$. We can think of the decoding problem in terms of the graph $\mathcal{G}$ as follows: The set of faults $F$ that occur specifies a set of fault nodes in $\mathcal{G}$. The syndrome $\sigma(F)$ is the set of check nodes of $\mathcal{G}$ that are connected to an odd number of fault nodes in $F$. A correction $\hat{F}$ is a set of fault nodes such that each node in $\sigma(F)$ is connected to an odd number of fault nodes in $\hat{F}$.
We interchangeably use bitstring and set notation in a way that is standard in the QEC literature, using the same conventions as in Ref. [Ott2025]. We use $j=1,2,\dots,N$ to index fault nodes. We then use $F$ to denote a set of fault nodes in two equivalent representations: as a set $F=\{j_{1},j_{2},\dots\}$, and as a bitstring $F\in\mathbb{F}_{2}^{N}$ with $F_{j}=1$ if, and only if, $j\in\{j_{1},j_{2},\dots\}$. Similarly, we use $\sigma$ to denote a set of check nodes both as a set $\sigma=\{i_{1},i_{2},\dots\}$ and as a bitstring $\sigma\in\mathbb{F}_{2}^{M}$ with $\sigma_{i}=1$ if, and only if, $i\in\{i_{1},i_{2},\dots\}$. When $U$ and $V$ are interpreted as sets, we use $U+V$ to denote their symmetric difference, consistent with $U+V$ being addition modulo two when $U$ and $V$ are interpreted as bitstrings. When $U$ and $V$ are interpreted as bitstrings, we use $U\cup V$ to denote the (non-exclusive) OR, consistent with $U\cup V$ being the union when $U$ and $V$ are interpreted as sets. Lastly, for $U$ and $V$ interpreted as sets, when $U\subseteq V$ we sometimes write VUV-U or $V\setminus U$ to specify set difference (which is also equivalent to $U+V$ since arithmetic is over $\mathbb{F}_{2}$). The syndrome $\sigma$ of a fault set $F=\{j_{1},j_{2},\dots\}$ is $\sigma=\mathcal{N}(j_{1})+\mathcal{N}(j_{2})+\dots=HF$, where arithmetic is over $\mathbb{F}_{2}$.
Figure 4: Tanner graphs for a simple noise model where the circuit consists of the measurement of each check of the code (toric code on the left and gross code on the right), and the set of faults consists only of an $X$ error on each qubit prior to the check measurements. Note that for the case of the gross code, we only draw the long-range edges (red) from one check, but all checks have similar edges which can be obtained by translation of those shown on the Torus.
QEC circuit decoding examples.
We benchmark our decoders on circuits for the $d=12$ (aka “gross”) and $d=18$ (aka “two-gross”) bivariate-bicycle codes of Refs. [bravyi2024high, bravyi2024github], using the same setup. These decoding examples consist of $d$ noisy QEC rounds (implemented with an explicit circuit that measures all code checks) followed by one noiseless round to project to the codespace and test for a logical error.
We assume the standard circuit-level noise model: preparations and measurements fail with probability $p$ (state flip or outcome flip, respectively); idle gates experience $X$, $Y$, or $Z$ with probability $p/3$ each; CNOTs suffer one of the $15$ nontrivial two-qubit Pauli errors with probability $p/15$ each. During each round, each check of the code is measured by the noisy circuit: we form a detector from the parity of a check’s consecutive measurement (corresponding to a row of $H$), and each distinct failure event corresponds to a column of $H$. The probability $p_{j}$ is set to the value of the probability of fault $j$ occuring (for example, $p_{j}=p/15$ if fault $j$ is an $XY$ error on a CNOT). We set $H_{ij}=1$ if fault $j$ flips detector $i$, and set $H_{ij}=0$ otherwise.
Each row of $A$ corresponds to a logical operator generator (and as such there are $2k$ rows for a code encoding $k$ logical qubits). Each fault produces a residual Pauli error on the data qubits, and we set $A_{ij}=1$ if the residual Pauli of fault $j$ anticommutes with logical operator generator $i$, and set $A_{ij}=0$ otherwise. Because these codes are CSS [calderbank1996good, steane1996multiple], $X$ and $Z$ decoding can be decoupled; here we consider $Z$-type decoding, with $X$-type decoding analogous. This is done by separating rows of the decoding matrix $H$ into detectors formed from $X$-type check outcomes, and retaining only columns with support on those rows. The same set of columns are used to select a submatrix of $A$ and a subvector of $\mathbf{p}$. The resulting $Z$-type decoding problem is represented by a binary decoding matrix $H_{Z}\in\mathbb{F}_{2}^{M\times N}$, a binary logical action matrix $A_{Z}\in\mathbb{F}_{2}^{K\times N}$, and a fault probability vector $\mathbf{p}_{Z}\in[0,1]^{N}$. The decoding matrices that arise in our two examples are $936\times 8784$ and $2736\times 26208$ for the gross and two-gross codes respectively.
2.2 Real-time decoding and latency-tail conditions
A large-scale quantum computation may involve a sequence of more than $10^{10}$ QEC cycles which implement a sequence of fault-tolerant operations (‘operations’ for short). Here we discuss constraints that arise to ensure decoding keeps pace with the continuous stream of syndrome data produced by this sequence of QEC cycles.
In practice, decoding of the stream of syndrome data is broken into many smaller decoding problems, each involving a small overlapping batch of $\sim d$ consecutive cycles supported on a small number of code blocks. For presentation clarity, we make the simplifying assumption that each decoding problem is disjoint: it corresponds to a single operation consisting of a contiguous sequence of QEC cycles on one code block, and different operations consist of non-overlapping sequences of cycles. This assumption lets us present the core ideas of real-time decoding without the added complexity of decoding problems that span overlapping windows of cycles and code blocks.
Computational slowdown from decoding latency.
Many fault-tolerant architectures, including surface-code [fowler2012surface] and qLDPC-based [yoder2025tour] schemes, implement non-Clifford operations via state injection. This introduces branching points in the computation, where the next operation depends on the decoder output from a previous operation (see Figure LABEL:fig:branching-and-latency). If the decoder requires $T_{\text{dec}}$ QEC cycles to return a result (for now we assume the decoder always takes the same amount of time for every decoding instance), the computation cannot proceed past the branch until that time.